alones (alones):

GEEEZS oboly plz show me how to check my answer for 5y = x - 3.

Vocaloid (vocaloid):

what is your answer?

alones (alones):

umm i got 5

Vocaloid (vocaloid):

you should have two numbers, one for x and one for y

Vocaloid (vocaloid):

is there more information given?

alones (alones):

yeah here_Find the slope of the line whose equation is 5y = x - 3.

Vocaloid (vocaloid):

oh, ok, that makes more sense

Vocaloid (vocaloid):

start by dividing both sides by 5

alones (alones):

okayi i'll will

Vocaloid (vocaloid):

let me know what you get

alones (alones):

i got 0

Vocaloid (vocaloid):

uh not quite...

Vocaloid (vocaloid):

what is 5y/5?

alones (alones):

1and y

alones (alones):

@AlexandervonHumboldt2 can you plz help me i'm sitting here and waisting my time..

Vocaloid (vocaloid):

ack, sorry, now divide the other side by 5

OpenStudy (alexandervonhumboldt2):

ok so: 5y = x - 3. just divide both sides by 5: 5y/5=(x-3)/5 y=1/5x-3/5 in y=mx+b form m is slope.

alones (alones):

ugh i did and i got 0

OpenStudy (alexandervonhumboldt2):

0? how?

alones (alones):

O_O i mean 1 and -2

OpenStudy (alexandervonhumboldt2):

m=1/5

OpenStudy (alexandervonhumboldt2):

m is slope

alones (alones):

O_o I got it so u find the slope and it becomes 1/5 is that how it suppose to be??